# Machine Unlearning in Hyperbolic vs. Euclidean Multimodal Contrastive Learning: Adapting Alignment Calibration to MERU

### v1

@@ 4.1 Experimental Setup
To investigate the impact of unlearning class-concepts from contrastive representation spaces, we focus on selectively removing the semantic relationship between image and text concepts, such as dogs, cats, food, and plants. Furthermore, we compare the efficacy of unlearning in both Euclidean and hyperbolic latent spaces with CLIP and MERU respectively.
CLIP and MERU [^1] Models availabe in MERU’s GitHub repository with ViT-S as image and text encoders [3], pretrained on the RedCaps dataset, are used as the orginal models.
During experiments only a subset of 7M image-text pairs from the original RedCaps dataset was accessible. This, referred to as RedCaps2, serves as the primary data source. Instances related to dogs, cats, food, and plants are categorized into four distinct high-level semantic concepts. These categorizations inform three ablation studies designed to evaluate unlearning capabilities 4.3.
The unlearning procedure operates on balanced mini-batches with $N$ image-text pairs from the retain set and $N$ pairs from the unlearn set, resulting in a batch size of $2N$, with $N=160$. This ensures fair representation of both retained and forgotten training concepts.
@@ 4.2 Evaluation
4.2.1 Zero-shot image classification
@@
4.2.2 Latent Space Visualizations
Image and text embeddings are visualized using both T-SNE [26] and hyperbolic T-SNE [11]. These visualizations may reveal structural changes in the representation space after unlearning. We look particularly if there is a distortion in the alignment between the image and text embedding of the concepts to forget.
4.2.3 Datasets
@@ 4.3 Ablation Studies
4.3.1 Alignment Calibration Components
This ablation assesses the impact of different components in the AC on unlearning performance. The weights from Eq. 8 ($\alpha,\beta,\gamma\in[0,1]$), are systematically varied to determine the optimal configuration for unlearning a single concept (dogs) in both CLIP and MERU, based on a zero-shot classification evaluation on CIFAR-10. Additionally, the overall weight of the AC components is also tested ($\varepsilon\in[0,0.1]$). As baselines, we evaluate also the orignal CLIP and MERU (o-C and o-M), both models finetuned only with the retain dataset (f-C-R and f-M-R), and both models finetuned on the entire dataset (f-C and f-M).
4.3.2 Hyperbolic-Specific Parameters
This ablation tests parameters unique to HAC, including entailment weights ($\omega_{r},\omega_{u}\in[0,1]$) and hyperbolic norm regularization ($\lambda\in[0,2]$). The analysis explores how these hyperbolic-specific factors influence unlearning, also evaluating on zero-shot classification for CIFAR-10.
4.3.3 Scaling Analysis
By selecting the best parameters based on the previous ablations, we perform a full evaluation on diverse evaluation datasets on three experiments. Experiment A, consists of removing the concept-class dogs alone. Removing dogs serves as an ideal initial test case since it presents particular challenges for unlearning, since the model can easily mistake a dog for a similar mammal (e.g., horses, cats). Experiment B extends this by removing both dogs and cats. Testing whether the unlearning mechanism can handle multiple related concepts simultaneously, as dogs and cats share numerous visual and semantic features. Finally, Experiment C consists of unlearning dogs, cats, food, and plants, to test how each geometry handles large-scale concept removal without catastrophic forgetting of non-targeted classes.
@@ 5.1 Alignment Calibration Components
We begin our analysis by examining how varying the key components of AC affects unlearning performance. Table 1 presents the effects of varying positive alignment calibration ($\beta$). For better numerical stability of the experiments, a small regularization norm with $\lambda=0.1$.
Table 1: Zero-shot classification accuracy in retain and forget sets, varying positive alignment calibration. Best trade-off performance in green. Best specific in bold.
@@
$\alpha,\gamma$ $\beta$ R-acc F-acc R-acc F-acc
@@
Our findings reveal several key insights about the role of $\beta$. First, increasing $\beta$ leads to a significant reduction in forget accuracy (F-acc) in both AC and HAC, confirming that this component effectively pushes positive image-text pairs further apart in their respective embedding spaces. This is consistent with previous research showing that controlled modification of alignment between positive pairs can lead to selective forgetting [29, 27].
However, we observe a dramatic difference in unlearning effectiveness. In HAC, introducing even a small $\beta=0.25$, it causes forget accuracy to plummet on CIFAR-10. This suggests that hyperbolic geometry amplifies the effect of positive alignment calibration, potentially due to its natural exponential expansion. The cost of this enhanced forgetting capability is reflected in retain accuracy (R-acc). HAC experiences a more substantial drop in retain performance compared to AC when $\beta$ is increased. Despite this trade-off, the relative difference between the drop in retain accuracy versus forget accuracy is more favorable in hyperbolic space, indicating superior unlearning efficiency.
On the other side, table 2 examines the effects of varying negative alignment calibration ($\alpha$) and performance preserving ($\gamma$) weights while maintaining a fixed positive alignment calibration ($\beta=0.5$).
Table 2: Zero-shot classification accuracy in retain forget sets, varying negative alignment calibration and performance preserving. Best trad-off performance in green. Best specific performance in bold.
@@
$\alpha,\gamma$ $\beta$ R-acc F-acc R-acc F-acc
@@
For AC, we observe that the retain accuracy (R-acc) remains relatively stable across different $\alpha$ and $\gamma$ values, while forget accuracy (F-acc) shows more substantial fluctuations. This suggests that in Euclidean space, modifying $\alpha$ and $\gamma$ has a more pronounced effect on the model’s ability forget the selected concept. In stark contrast, HAC demonstrates remarkably consistent forgetting capabilities across all parameter configurations, maintaining near-zero forget accuracy regardless of $\alpha$ and $\gamma$ values. This remarkable stability in forgetting performance suggests that in hyperbolic space, once the positive alignment calibration establishes a forgetting directive, it dominates the unlearning process to such an extent that adjustments to other components have minimal on forget accuracy. However, the retain accuracy in HAC shows greater sensitivity to parameter changes.
These findings suggest fundamentally different optimization landscapes between Euclidean and hyperbolic spaces for concept unlearning. In Euclidean geometry, the three components of alignment calibration work in relative harmony, requiring careful balancing to achieve both retention and forgetting. In hyperbolic geometry, the positive alignment calibration component exerts such a dominant influence that it essentially decouples the forgetting process from the other components.
@@ 5.2 Hyperbolic-Specific Parameters
Having established the fundamental differences between AC and HAC, we now explore parameters that are unique to the hyperbolic geometry. Table 3 illustrates the impact of varying entailment loss weights on HAC performance.
Table 3: Zero-shot classification accuracy in retain and forget sets, varying the weight entailment losses. Best trade-off performance in green. Best specific performance in bold.
@@
$\epsilon$ $\omega_{r}$ $\omega_{f}$ R-acc F-acc R-acc F-acc
@@
Our analysis demonstrates that entailment losses play a pivotal role in hyperbolic unlearning quality. When AC components are deactivated ($\varepsilon=0$), we observe not only a failure to achieve unlearning but actually an increase in forget accuracy on CIFAR-10. This finding suggests that entailment losses alone may inadvertently strengthen the association between text and image embeddings of the target concept, increasing their relative similarity, instead of decreasing it.
The introduction of AC components ($\varepsilon>0$) dramatically changes this dynamic. With $\varepsilon=0.05$ or $0.1$, we achieve complete forgetting (0% F-acc) on CIFAR-10 when prioritizing the forget entailment ($\omega_{f}=1.0$). This demonstrates the synergistic effect between alignment calibration and entailment losses in hyperbolic space. Similar patterns emerge for O-IIIT Pets, yielding near-perfect forgetting (0.08% F-acc) while maintaining reasonable retain accuracy (60.7%).
The balance between retain and forget entailment weights ($\omega_{r}$ and $\omega_{f}$) proves crucial for optimal unlearning. When $\omega_{r}>\omega_{f}$, we observe significantly higher forget accuracy across all configurations, indicating incomplete concept removal. This supports the theoretical understanding that in hyperbolic space, the precise positioning of embeddings within semantic cones strongly influences model behavior [7].
Finally, the hyperbolic norm regularization term also has remarkable implications on the quality of concept removal. Beyond providing better stability, the regularization term actively improves unlearning quality, as shown in Table 4. With $\lambda=0.5$, we achieve optimal performance across both datasets, completely eliminating forget accuracy on CIFAR-10 while improving retain accuracy. This regularization effect is particularly beneficial for O-IIIT Pets, where it reduces forget accuracy while maintaining strong retain performance. The mechanism behind this improvement involves pulling text embeddings for the forgotten concept toward the origin of the hyperboloid, causing other images that were previously misclassified, to be reclassified under other categories, provably their category, increasing retain accuracy. This is especially important for fine-grained classification in O-IIIT Pets, where the regularization helps disentangle dogs from similar categories like cats.
Table 4: Zero-shot classification accuracy in retain and forget sets, varying the hyperbolic norm regularization. Best trade-off performance in green. Best specific performance in bold.
@@
$\lambda$ R-acc F-acc R-acc F-acc
@@
These results highlight the fundamental geometric differences in concept unlearning between Euclidean and hyperbolic spaces, with hyperbolic geometry enabling more definitive concept removal at the cost of higher sensitivity to parameter tuning.
@@ 5.3 Scaling Analysis
Having established the optimal configurations for both AC and HAC, we now investigate how these approaches scale when progressively increasing the complexity of the unlearning task. Table 7 presents a comprehensive comparison across multiple datasets when unlearning: (A) only ”dog”; (B) ”dog” and ”cat”; and (C) ”dog”, ”cat”, ”food”, and ”plants”.
Table 5: Zero-shot classification accuracy in retain set (R-acc) and forget set (F-acc), across different tasks, after unlearning: (A) ”dog”; (B) ”dog” and ”cat”; (C) ”dog”, ”cat”, ”food” and ”plants”. We report results for both CLIP and MERU after alignment calibration using the optimal configuration from Section 4.3.1. Values in bold indicate better at retaining or unlearning across A, B and C.
@@
We observe a clear trade-off pattern between retain accuracy and forget accuracy across all experiments. AC consistently maintains higher retain accuracy across all datasets and concept sets compared to HAC. However, this advantage comes at a significant cost: AC’s forget accuracy increases dramatically as the unlearning task scales, indicating incomplete concept removal.
In stark contrast, HAC demonstrates superior forgetting capabilities across all scaling scenarios. Even when tasked with unlearning four diverse concepts simultaneously (experiment C), HAC maintains remarkably low forget accuracies. Suggesting that HAC’s leverages the hierarchical hyperbolic structure to effectively remove multiple concepts from the embedding space without cross-interference.
These scaling experiments reveal that HAC offers more robust concept removal at scale, while AC better preserves performance on retained concepts. This fundamental trade-off highlights the complementary strengths of each geometric approach and suggests that the choice between them should be guided by whether the priority is complete concept removal or minimal disruption to retained knowledge.
@@ 6 Latent Space Visualizations
Figure 1: Latent space visualizations with T-SNE of CLIP and MERU before and after removing the concept-class ”dog”. $\triangle$ refer to text embeddings, $\circ$ to image embeddings, and colors to dogs cats, horses, and buses.
@@
Figure 2: Latent space visualizations with hyperbolic T-SNE of MERU before and after removing the concept-class ”dog”. $\triangle$ refer to text embeddings, $\circ$ to image embeddings, and colors to dogs cats, horses, and buses.
The T-SNE visualizations in Figure 1 reveal several key patterns. First, both AC and HAC induce a general separation between image and text representations after unlearning, indicating that cross-modal alignment disruption is a common mechanism regardless of geometry. However, the degree and nature of this separation differ markedly between models. In the Euclidean space of CLIP (Figures 1a-b), both text and image embeddings maintain their class-wise clustering structure after unlearning, with dog representations remaining distinct but disconnected from each other across modalities. In contrast, MERU’s hyperbolic space (Figures 1c-d) exhibits a more profound reorganization of semantic relationships. While image embeddings preserve their class separation, text embeddings show significant restructuring, with dog text representations (pink triangles) migrating toward the region occupied by cat text embeddings (red triangles). This visualization corroborates our quantitative findings that hyperbolic geometry facilitates more complete concept removal by leveraging semantic hierarchies—in this case, potentially positioning ”dog” as a subconcept of a more general ”animal” category that includes cats. However, this poses a potential limitation of hyperbolic unlearning: while effective at repositioning target concepts, the structural interdependence of the hierarchical space means that semantically adjacent concepts (like cats when removing dogs) experience collateral disruption in their text embeddings, potentially explaining the lower retention accuracy observed in HAC compared to AC.
The hyperbolic T-SNE visualizations in Figure 2 provide additional insights into MERU’s unlearning dynamics. Following unlearning, we observe a dramatic expansion of the embedding space, with image representations pushed substantially farther from the origin while text embeddings remain relatively close to it. This pattern aligns with hyperbolic geometry’s exponential expansion property and illustrates how HAC exploits this characteristic. Particularly notable is the positioning of dog text embeddings near the origin after unlearning, effectively placing them ”behind” cat text embeddings from the perspective of dog images. This spatial reorganization explains the superior unlearning results evaluated in zero-shot classification, dog images now find stronger alignment with text embeddings of other classes because their original text concept has been repositioned closer to the origin within the semantic hierarchy.

### v2

@@ 4.1 Experimental Setup
To investigate the impact of removing concepts from contrastive representation spaces, we focus on selectively disrupting the semantic relationship between image and text of concepts, such as dogs, cats, food, and plants. Both CLIP and MERU models have ViT-S [5] as image encoder and a 12-layer, 512 wide Trasnformer as text encoder [30]. They are pretrained on the RedCaps dataset and used as the orginal models [^1] Models availabe in MERU’s GitHub repository that require unlearning.
During experiments only a subset of 7M image-text pairs from the original RedCaps dataset was accessible. This, referred to as RedCaps2, serves as the original dataset. Instances related to dogs, cats, food, and plants are categorized into four distinct high-level semantic concepts, defining the different forget sets, see Sec. 11 in supp. material for more details on the set up forget sets. The unlearning procedure operates on balanced mini-batches with $N$ image-text pairs from the retain set and $N$ pairs from the unlearn set, resulting in a batch size of $2N$, with $N=160$.
@@ 4.2 Evaluation
Zero-shot image classification
@@
Latent Space Visualizations
Image and text embeddings are visualized using both T-SNE [35] and hyperbolic T-SNE [13]. The latter is an extension of T-SNE which uses hyperbolic student’s t-distribution to capture the hierarchical relationships between embeddings. These visualizations may reveal structural changes in the representation space after unlearning. We look particularly if there is a distortion in the alignment between the image and text embedding of the concepts to forget.
Datasets
@@ 4.3 Ablation Studies
Alignment Calibration Components
We assess the impact of the different alignment components on unlearning performance. We vary the weights from Eq. 11 ($\alpha,\beta,\gamma\in[0,1]$), unlearning a single concept (dogs) in both CLIP and MERU. As baselines, we consider the orignal CLIP and MERU (o-C and o-M), both models finetuned only with the retain dataset (f-C-R and f-M-R), and both models finetuned on the entire dataset (f-C and f-M).
Hyperbolic-Specific Parameters
We ablate parameters unique to HAC: entailment weights ($\omega_{r},\omega_{u}\in[0,1]$) and hyperbolic norm regularization ($\lambda\in[0,2]$). We analyze how these hyperbolic-specific factors influence unlearning, evaluating on zero-shot classification for CIFAR-10.
Scaling Analysis
We evaluate on diverse datasets and three experiments. Experiment A, consists of removing the concept-class dogs alone. Removing dogs serves as an ideal initial test case, since it has a strong semantic relationship with other concepts to retain (e.g., horses, cats). Experiment B extends this by removing both dogs and cats. Testing whether the unlearning mechanism can handle multiple related concepts simultaneously. Finally, Experiment C consists of unlearning dogs, cats, food, and plants, to test how each geometry handles large-scale concept removal without catastrophic forgetting of non-targeted concepts.
@@ 5.1 Alignment Calibration Components
We begin our analysis by examining how changing the key components of AC affects unlearning performance. Table 1 presents the effects of varying positive alignment calibration ($\beta$). For better numerical stability of the experiments, we introduce a small regularization norm with $\lambda=0.1$ for HAC. We observe that increasing $\beta$ leads to a significant reduction in forget accuracy (F-acc) in both AC and HAC, confirming that this component effectively pushes positive image-text pairs further apart in their respective embedding spaces and leading to zero-shot misclassifications. This is consistent with previous research showing that controlled modification of alignment between positive pairs can lead to selective forgetting [38, 36]. However, we observe a dramatic difference in unlearning effectiveness. In HAC, introducing even a small $\beta$, it causes forget accuracy to plummet on CIFAR-10. This suggests that hyperbolic geometry amplifies the effect of positive alignment calibration, potentially due to its natural exponential expansion. The cost of this enhanced forgetting capability is reflected in retain accuracy (R-acc). HAC experiences a more substantial drop in retain performance compared to AC when $\beta$ is increased. Despite this trade-off, the relative difference between the drop in retain accuracy versus forget accuracy is more favorable in hyperbolic space, indicating superior unlearning efficiency.
Table 1: Zero-shot classification accuracy in retain and forget sets, varying positive alignment calibration. Largest difference in retain and forget performance in green. Best value for each column and geometry in bold.
@@
$\alpha,\gamma$ $\beta$ R-$\uparrow$ F-$\downarrow$ R-$\uparrow$ F-$\downarrow$
@@
On the other side, table 2 examines the effects of varying negative alignment calibration ($\alpha$) and performance preserving ($\gamma$) weights while maintaining a fixed positive alignment calibration. For AC, we observe that the retain accuracy (R-acc) remains relatively stable across different $\alpha$ and $\gamma$ values, while forget accuracy (F-acc) shows more substantial fluctuations. This suggests that in Euclidean space, modifying $\alpha$ and $\gamma$ has a more pronounced effect on the model’s ability forget the selected concept. In contrast, HAC demonstrates remarkably consistent forgetting capabilities across all parameter configurations, maintaining near-zero forget accuracy regardless of $\alpha$ and $\gamma$ values, while dampening retain accuracy. This remarkable behavior suggests that in hyperbolic space, once the positive alignment calibration establishes a forgetting directive, it dominates the unlearning process to such an extent that adjustments to other components extend the disruption to the retain set.
Table 2: Zero-shot classification accuracy in retain forget sets, varying negative alignment calibration and performance preserving. Largest difference in retain and forget performance in green. Best value for each column and geometry in bold.
@@
$\alpha,\gamma$ $\beta$ R-$\uparrow$ F-$\downarrow$ R-$\uparrow$ F-$\downarrow$
@@ 5.2 Hyperbolic-Specific Parameters
We now explore parameters that are unique to the hyperbolic geometry. Table 3 illustrates the impact of varying entailment loss weights on HAC performance. When AC components are deactivated ($\varepsilon=0$), we observe not only a failure to achieve unlearning but actually an increase in forget accuracy on CIFAR-10. This finding suggests that entailment losses alone may inadvertently strengthen the association between text and image embeddings of the target concept, increasing their relative similarity, instead of decreasing it.
Table 3: Zero-shot classification accuracy in retain and forget sets, varying the weight entailment losses. Largest difference in retain and forget performance in green. Best value for each column and geometry in bold.
@@
$\epsilon$ $\omega_{r}$ $\omega_{f}$ R-$\uparrow$ F-$\downarrow$ R-$\uparrow$ F-$\downarrow$
@@
The introduction of AC components ($\varepsilon>0$) dramatically changes this dynamic. Furthermore, when prioritizing the forget entailment ($\omega_{f}=1.0$) complete forgetting (0% F-acc) on CIFAR-10 is achieved. This demonstrates the synergistic effect between alignment calibration and entailment losses in hyperbolic space. Similar patterns emerge for O-IIIT Pets, yielding near-perfect forgetting while maintaining reasonable retain accuracy.
The balance between retain and forget entailment weights ($\omega_{r}$ and $\omega_{f}$) proves crucial for optimal unlearning. When $\omega_{r}>\omega_{f}$, we observe significantly higher forget accuracy across all configurations, indicating incomplete concept removal. This supports the theoretical understanding that in hyperbolic space, the precise positioning of embeddings within entailment cones strongly influences model behavior [9].
Finally, beyond providing better stability, the regularization term actively improves unlearning quality, as shown in Table 4. With $\lambda=0.5$, we largest retain-forget trade-off across both datasets. This regularization effect is particularly beneficial for retaining accuracy illustrated with O-IIIT Pets, where at the cost of increasing forgetting accuracy, we observe a strong retain performance. The mechanism behind this improvement involves pulling text embeddings for the forgotten concept toward the origin of the hyperboloid, causing other images that were previously misclassified, to be reclassified under other categories, provably their category, increasing retain accuracy. This is especially important for fine-grained classification in O-IIIT Pets, where the regularization helps disentangle dogs from similar categories like cats.
Table 4: Zero-shot classification accuracy in retain and forget sets, varying the hyperbolic norm regularization. Largest difference in retain and forget performance in green. Best value for each column and geometry in bold.
@@
$\lambda$ R-$\uparrow$ F-$\downarrow$ R-$\uparrow$ F-$\downarrow$
@@ 5.3 Scaling Analysis
Fixing hyperparametes to the combination found with largest forget-retain difference, for both AC and HAC, we now investigate how these approaches scale when progressively increasing the complexity of the unlearning task. Table 7 presents a comprehensive comparison across multiple datasets when unlearning: (A) only ”dog”; (B) ”dog” and ”cat”; and (C) ”dog”, ”cat”, ”food”, and ”plants”.
Table 5: Zero-shot classification accuracy in retain set (R-acc) and forget set (F-acc), across different tasks, after unlearning: (A) ”dog”; (B) ”dog” and ”cat”; (C) ”dog”, ”cat”, ”food” and ”plant”. We report results for both CLIP and MERU after alignment calibration using the optimal configuration from Section 4.3. Values in bold indicate better at retaining or unlearning across A, B and C. A blank space - indicate that for that experiment and dataset there is no forget or retain set.
Task Method Unlearn Set CIFAR-10[18] CIFAR-100[19] STL-10[1] O-IIIT Pets[26] Food101[3] Flowers102[25]
R-$\uparrow$ F-$\downarrow$ R-$\uparrow$ F-$\downarrow$ R-$\uparrow$ F-$\downarrow$ R-$\uparrow$ F-$\downarrow$ R-$\uparrow$ F-$\downarrow$ R-$\uparrow$ F-$\downarrow$
@@
We observe a clear trade-off pattern between retain accuracy and forget accuracy across all experiments. AC consistently maintains higher retain accuracy across all datasets and concept sets compared to HAC. However, this advantage comes at a significant cost: AC’s forget accuracy increases dramatically as the unlearning task scales, indicating incomplete concept removal and lower scaling capabilities. In stark contrast, HAC demonstrates superior forgetting capabilities across all scaling scenarios. Even when tasked with unlearning four diverse concepts simultaneously (experiment C), HAC maintains remarkably low forget accuracies. Suggesting that HAC’s leverages the hierarchical hyperbolic structure to effectively remove multiple concepts from the embedding space without cross-interference.
These experiments reveal that HAC offers more robust concept removal at scale, while AC better preserves performance on retained concepts. This fundamental trade-off highlights the complementary strengths of each geometric approach and suggests that the choice between them should be guided by whether the priority is complete concept removal or minimal disruption to retained knowledge. Section 12 in supplementary material complements our findings showcasing the confusion matrices foer CIFAR-10.
@@ 6 Latent Space Visualizations
Figure 1: Latent space visualizations with T-SNE of CLIP and MERU before and after removing the concept ”dog”. $\triangle$ refer to text embeddings, $\circ$ to image embeddings, and colors to dogs, cats, pizzas, and buses.
@@
Figure 2: Latent space visualizations with hyperbolic T-SNE of MERU before and after removing the concept ”dog”. $\triangle$ refer to text embeddings, $\circ$ to image embeddings, and colors to dogs, cats, pizzas, and buses.
The T-SNE visualizations in Figure 1 reveal several key patterns. First, both AC and HAC induce a general separation between image and text representations after unlearning, indicating that enlarging the modality gap is a common consequence of unlearning regardless of geometry. However, the degree and nature of this separation differ between models. In the Euclidean space of CLIP (Figures 1a-b), both text and image embeddings maintain their class-wise clustering structure after unlearning. In contrast, MERU’s hyperbolic space (Figures 1c-d) exhibits a more profound reorganization of semantic relationships. While image embeddings preserve their class separation, text embeddings show significant restructuring, with dog text representations (pink triangles) migrating toward the region occupied by cat text embeddings (red triangles). This visualization corroborates our quantitative findings that hyperbolic geometry facilitates complete concept removal by leveraging semantic hierarchies—in this case, potentially positioning ”dog” as a subconcept of a more general ”animal” category that includes cats. However, this poses a potential limitation of hyperbolic unlearning: while effective at repositioning target concepts, the structural interdependence of the hierarchical space means that semantically adjacent concepts (like cats when removing dogs) experience collateral disruption in their text embeddings, potentially explaining the lower retention accuracy observed in HAC compared to AC.
The hyperbolic T-SNE visualizations in Figure 2 provide additional insights into MERU’s unlearning dynamics. Following unlearning, we observe a dramatic expansion of the embedding space, with image representations pushed substantially farther from the origin while text embeddings remain relatively close to it. This pattern aligns with hyperbolic geometry’s exponential expansion property and illustrates how HAC exploits this characteristic. Particularly notable is the positioning of dog text embeddings near the origin after unlearning, effectively placing them ”behind” cat text embeddings from the perspective of dog images. This spatial reorganization explains the superior unlearning results evaluated in zero-shot classification, dog images now find stronger alignment with text embeddings of other classes because their original text concept has been repositioned closer to the origin within the semantic hierarchy. Complementary illustrations can be found in Section 14 in supplementary material.